OpenStudy (anonymous):

Subtract. Simplify, if possible. (2-v)/(v-6)-(5V-4)/(6-V)

OpenStudy (anonymous):

\[\frac{2-v}{v-6} - (\frac{5v-4}{6-v} \times \frac{-1}{-1}) \]\[=\frac{2-v-4+5v}{v-6} = \frac{4v-2}{v-6}\]